OpenStudy (anonymous):

Please help solve y"-2(y')^2=0 using dy/dx=p

OpenStudy (anonymous):

\[y"+2(y')^{2}=0 using \frac{ dy }{ dx }=\rho\]

OpenStudy (experimentx):

let dy/dx = u then you get du/dx + 2u^2 = 0 <-- this is first order DE use separation of variables du/u^2 + 2dx = 0 -1/u + 2x + c_1 = 0 u = 1/(2x + c_1) again we have u = dy/dx dy/dx = 1/(2x+c_1) dy = dx/(2x+c_1) integrate both sides you will get y = 1/2 log(2x+c_1) + c_2

OpenStudy (anonymous):

The quest asks if u can use dy/dx=p

OpenStudy (anonymous):

@ ajprincess

OpenStudy (anonymous):

@moyo30 experimentX is right in solving this ques. He just used dy/dx = u instead of dy/dx = p .

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i c it now